# Optimal Non-Adaptive Group Testing with One-Sided Error Guarantees

### v3

@@ Abstract
The group testing problem consists of determining a sparse subset of defective items from within a larger set of items via a series of tests, where each test outcome indicates whether at least one defective item is included in the test. We study the approximate recovery setting, where the recovery criterion of the defective set is relaxed to allow a small number of items to be misclassified. In particular, we consider one-sided approximate recovery criteria, where we allow either only false negative or only false positive misclassifications. Under false negatives only (i.e., finding a subset of defectives), we show that there exists an algorithm matching the optimal threshold of two-sided approximate recovery. Under false positives only (i.e., finding a superset of the defectives), we provide a converse bound showing that the better of two existing algorithms is optimal.
@@ I Introduction
There are several different flavors of group testing, each having different results on the minimum number of tests needed for reliable recovery. For instance, one can consider both adaptive testing where previous test results can be used to inform one’s decisions in designing subsequent tests, and non-adaptive testing where all tests are designed before any testing takes place. Additionally, one can consider both the linear regime where $k=\Theta(n)$ (i.e. $k$ scales linearly with $n$) or the sparse regime where $k=\Theta(n^{\theta})$ for some $\theta\in\left(0,1\right)$ (i.e. $k$ scales sub-linearly relative to $n$).
@@
  • In small-error probabilistic group testing with exact recovery, we require that $\mathbb{P}(\hat{S}\neq S)\to 0$ as $n\to\infty$, where the randomness is with respect to $S$ and/or the possibly randomized test design and decoding algorithm. This relaxation significantly reduces the required number of tests to $O(k\log n)$.
  • One can further relax small-error probabilistic group testing to the approximate recovery criterion, in which instead of requiring $\hat{S}=S$ exactly, we allow for a small number of false positives and false negatives in the estimate. While this typically does not improve the scaling of the number of tests, it can significantly improve the constant factors.
@@ I-A Problem Setup
Given a set of items $n]\coloneqq\{1,2,\dots n\}$, there exists some set $S\subseteq[n]$ of defective items with $\lvert S\rvert=k$, where $k=\Theta(n^{\theta})$ for some $\theta\in(0,1)$ and $k$ is known a priori. The set of all possible values of $S$ is denoted by $\mathcal{S}$}, and $S$ is assumed to be uniformly distributed over all $n\choose k$ elements in $\mathcal{S}$}, which is known as the combinatorial prior. We also consider in certain parts of our proofs the i.i.d prior (e.g., as an intermediate step), where each item is defective independently with some probability $q\in(0,1)$.
@@
where $\bigvee$ represents the ‘OR’ operator, $Y_{t}$ is the $t$-th entry of $\mathbf{Y}$, and $\mathbf{X}_{t}=({\mathsf{X}}_{t1},{\mathsf{X}}_{t2},\dots,{\mathsf{X}}_{tn})\in\{0,1\}^{n}$ is the $t$-th row of $\mathsf{X}$}. Given $\mathsf{X}$} and $\mathbf{Y}$, a decoder forms an estimate $\hat{S}$ of the defective set $S$.
@@
While previous works on approximate recovery focused mainly on this symmetric two-sided criterion, our focus in this paper is instead on to the following problems, which represent the extremes of placing unequal weighting on false positives and false negatives:
@@ I-B Prior Work
Exact recovery. Dorfman’s original scheme [1] consisted of splitting up the items into disjoint groups, and including all items of each group in a single test. All items in negative tests are marked as non-defective, while the remaining items in positive tests are subsequently tested individually. While this does indeed beat individual testing, such a scheme is highly suboptimal when the number of defectives is sufficiently sparse relative to the number of items.
@@
In the sparse setting (namely, $k=\Theta(n^{\theta})$ with $\theta\in(0,1)$), there exists a ubiquitous lower bound of $T\geq\big(k\log_{2}\frac{n}{k}\big)(1-o(1))$ on the number of tests required to obtain a vanishing probability of error, regardless of adaptivity or non-adaptivity. This bound is known as the counting bound [13]. The intuition behind this bound stems from the fact that $\log_{2}{n\choose k}=\big(k\log_{2}\frac{n}{k}\big)\big(1+o(1)\big)$ bits are required to describe the defective set, and each test can provide at most 1 bit of information since the test outcomes are binary. This intuition can easily be made precise using standard information theoretic tools such as Fano’s inequality [14, Theorem 1], which implies we cannot obtain an arbitrarily small probability of error with fewer tests. This result has also been strengthened to obtain a strong converse bound stating that $\mathbb{P}(\hat{S}\neq S)\to 1$ as $n\to\infty$ when $T\leq\left(k\log_{2}\frac{n}{k}\right)(1-\varepsilon)$ for arbitrarily small $\varepsilon>0$ [13, Thm. 3.1].
@@
The counting bound can be met when adaptive testing is used for all $\theta\in(0,1)$ using a generalized binary splitting algorithm due to Hwang [15]. The counting bound can also be achieved for $\theta\in\big(0,\frac{\log 2}{1+\log 2}\big)\approx(0,0.409)$ non-adaptively, while a converse bound specific to the non-adaptive setting shows that more tests are required for $\theta\geq\frac{\log 2}{1+\log 2}$ [16]. Furthermore, [16, Theorem 2] shows that there exists a polynomial-time algorithm matching this bound in the non-adaptive setting. (Prior to their work, the bound was showed to be information-theoretically achievable in [17].)
@@
Approximate recovery. There is also a converse bound specific to the approximate recovery setting [18, Theorem 3], which states that the high probability recovery of a set $\hat{S}$ such that $\max\big\{\lvert\hat{S}\setminus S\rvert,\lvert S\setminus\hat{S}\rvert\big\}\leq\beta k$ for fixed $\beta\in(0,1)$ is not possible when the number of tests satisfies $T\leq(1-\beta)\left(k\log_{2}\frac{n}{k}\right)(1-\varepsilon)$ for arbitrarily small $\varepsilon>0$. When an equal number of false positive and false negative items are allowed (i.e., two-sided approximate recovery), a matching upper bound can be achieved [12]. The polynomial-time algorithm from [16] can also be used to achieve the same bound.
@@ I-C Achievable Rates Adapted from Prior Works
With the definitions from Section I-A in mind, we now state the previously best known results for both problems, as well as strengthening such results slightly by allowing $\eta^{-}$ and $\eta^{+}$ in Definitions 2 and 3 to decay to zero at a certain (sufficiently slow) speed.
@@
Then if $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in[0,1)$, and $\eta^{+}=k^{o(1)}$, the rate $R^{*}$ is achievable in both the SUBSET and SUPERSET problems.
The condition $\eta^{+}=k^{o(1)}$ covers a range of possible values, including not only $\eta^{+}=\Theta(1)$ but also the case of growing sufficiently slowly (e.g., as $(\log k)^{c}$ for some $c>0$). Moreover, the condition on $\eta^{-}$ allows not only $\eta^{-}=\Theta(1)$ and “slow” decay to zero at rate $k^{-o(1)}$, but also faster (polynomial) decay; by letting $\lambda$ be close to $1$ we can allow $\eta^{-}=\big(\frac{k}{n}\big)^{1-\varepsilon}$ for arbitrarily small $\varepsilon>0$, amounting to at most $k\big(\frac{k}{n}\big)^{1-\varepsilon}$ false negatives in the SUBSET problem. Note also that the term $\zeta(\theta)$ is the optimal exact recovery rate (among all designs) from [16], and is present because achieving exact recovery implies solving both problems even with $\eta^{+},\eta^{-}=0$. This term is dominant for $\theta<\frac{1}{2}$, and for any such $\theta$ the quantity $\eta^{-}k=\Theta(k(\frac{k}{n})^{\lambda})$ is $o(1)$ for $\lambda$ sufficiently close to one, which means that the “approximate recovery” guarantee of SUBSET is in fact equivalent to exact recovery.
For the SUBSET problem, the rate $\log 2\approx 0.693$ is achievable using the Definite Defectives (DD) algorithm along with the near-constant tests-per-item design. See [27] for the algorithm studied under Bernoulli designs, [19] for its study under the improved design, and [8, Section 5.1] for an overview of the extension to approximate recovery.
For the SUPERSET problem, the rate of $\log 2$ is achievable using the Combinatorial Orthogonal Matching Pursuit (COMP) algorithm. See [14] for the algorithm and Bernoulli designs, [19] for its study under the improved test design, and [8, Section 5.1] for an overview of the extension to approximate recovery.
Since the details on these extensions are omitted from [8, Section 5.1], and since the outline therein only covers the case that $\eta^{-},\eta^{+}=\Theta(1)$, we detail the proof of Theorem 1 in Appendix -D.
@@ I-D Contributions
  • We provide a new algorithm for the SUBSET problem that achieves a rate of 1 for all $\theta\in(0,1)$ when $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in[0,\frac{1}{2})$, thus matching the counting bound and beating the best existing achievability result.
@@ I-E Notation
We generally use bold face letters to denote vectors, lowercase to denote scalars and capital letters to denote random variables (though the number of tests $T$ is an exception). We use standard Bachmann-Landau notation $o(\cdot),\ \omega(\cdot),\ O(\cdot),\ \Omega(\cdot),\ \Theta(\cdot)$ to denote asymptotic relations. For a matrix $\mathsf{M}$} (resp. vector $\mathbf{v}$) and set ${\mathcal{J}}\subseteq[n]$, ${\mathsf{M}}_{{\mathcal{J}}}$ denotes the sub-matrix of $\mathsf{M}$} (resp. sub-vector of $\mathbf{v}$) corresponding to the columns (resp. entries) indexed by $\mathcal{J}$}. If $\mathcal{J}$$\emptyset$ we use the convention that $\mathsf{M}$$_{\emptyset}$ (resp. $\mathbf{v}_{\emptyset}$) $\emptyset$. For a set $\mathcal{N}$} and some ${\mathcal{U}}\subseteq{\mathcal{N}}$, we denote ${\mathcal{U}}^{\mathsf{c}}\coloneqq{\mathcal{N}}\setminus{\mathcal{U}}$ as the complement of $\mathcal{U}$}. All logarithms have base $e$ unless specified otherwise (e.g., via $\log_{2}$).
@@ II Main Results
We can establish the tightness of these results by comparing them to existing results:
@@
Remark 1.
(Discussion on $\lambda$) While the range of $\lambda$ in Theorems 2 and 3 is not our main focus, we note that the range in Theorem 2 permits $\eta^{-}=\big(\sqrt{\frac{k}{n}}\big)^{1-\varepsilon}$ (i.e., $k\big(\sqrt{\frac{k}{n}}\big)^{1-\varepsilon}$ false negatives) for arbitrarily small $\varepsilon>0$, which is a stronger result than $\eta^{-}=\Theta(1)$ or $\eta^{-}=k^{-o(1)}$. Similarly, according to the range in Theorem 3, when $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$ the converse result holds even when $\eta^{+}=\big(\frac{n}{k^{2}}\big)^{1-\varepsilon}$ (i.e., $k\big(\frac{n}{k^{2}}\big)^{1-\varepsilon}\gg k$ false positives).
While we attempted to find the best possible range of $\lambda$ in Theorems 2 and 3 based on what our own analysis can give, we do not claim these to be the best possible among all designs. In particular, we expect that adopting a near constant column weight design for SUBSET would lead to higher range of $\lambda$ being allowed. [^1] One way of seeing this is to observe that setting $\theta<\frac{1}{3}$ and $\lambda\approx\frac{1}{2}$ in Theorem 2 permits $\eta^{-}<\frac{1}{k}$, which is equivalent to exact recovery. This is consistent with exact recovery being possible for all $\theta\leq\frac{1}{3}$ under the Bernoulli design. For the near-constant column weight design, this range increases to $\theta\leq\frac{\log 2}{1+\log 2}\approx 0.409$, which suggests that the range of allowed $\lambda$ should be higher, namely $\lambda\in[0,\log 2)$. We stick to the Bernoulli design since it simplifies the analysis, and because making $\lambda$ as high as possible is not our main focus.
Remark 2.
(Discussion on Computation) The optimal threshold for SUPERSET can be achieved with polynomial-time decoding. When the rate matches that of exact recovery, this follows directly from the results of [16] based on spatial coupling. The only other case is that in which the rate is $\log 2$, which is achieved by the Definitive Defectives (DD) algorithm (Theorem 1). For the SUBSET problem, however, our achievability proof (Theorem 2) is purely information-theoretic and uses a highly inefficient decoder. We leave it as an open problem as to whether the same can be achieved in polynomial time, possibly via the spatial coupling approach of [16].

### v4

@@ Abstract
The group testing problem consists of determining a sparse subset of defective items from within a larger set of items via a series of tests, where each test outcome indicates whether at least one defective item is included in the test. We study the approximate recovery setting, where the recovery criterion of the defective set is relaxed to allow some number of items (up to a certain specified threshold) to be misclassified. In particular, we consider one-sided approximate recovery criteria, where we allow either only false negative or only false positive misclassifications. Under false negatives only (i.e., finding a subset of defectives), we show that there exists an algorithm matching the optimal threshold of two-sided approximate recovery, albeit with exponential runtime. Under false positives only (i.e., finding a superset of the defectives), we provide a converse bound showing that the better of two existing algorithms is optimal.
@@ I Introduction
There are several different flavors of group testing, each having different results on the minimum number of tests needed for reliable recovery. For instance, one can consider both adaptive testing where previous test results can be used to inform one’s decisions in designing subsequent tests, and non-adaptive testing where all tests are designed before any testing takes place. Moreover, one can consider noiseless testing, where the correct outcome is always observed deterministically, or noisy testing, where some randomness is introduced to the test outcomes (e.g., test outcomes are flipped independently with some probability). Finally, one can consider both the linear regime where $k=\Theta(n)$ (i.e., $k$ scales linearly with $n$) or the sparse regime where $k=\Theta(n^{\theta})$ for some $\theta\in\left(0,1\right)$ (i.e., $k$ scales sub-linearly relative to nn). [^1] There is also a range of $k$ such that $k=n^{1-o(1)}$ but $k=o(n)$, as considered in [5], but this regime is narrow and less commonly considered compared to the other two discussed here.
@@
  • In small-error probabilistic group testing with exact recovery, we require that $\mathbb{P}(\hat{S}\neq S)\to 0$ as $n\to\infty$, where the randomness is with respect to $S$ (chosen uniformly at random) and/or the possibly randomized test design and decoding algorithm. This relaxation significantly reduces the required number of tests to $O(k\log n)$.
  • One can further relax small-error probabilistic group testing to the approximate recovery criterion, in which instead of requiring $\hat{S}=S$ exactly, we allow for a small number of false positives and/or false negatives in the estimate. While this typically does not improve the scaling of the number of tests, it can significantly improve the constant factors.
@@ I-A Problem Setup
Given a set of items $n]\coloneqq\{1,2,\dots,n\}$, there exists some set $S\subseteq[n]$ of defective items with $\lvert S\rvert=k$, where $k=\Theta(n^{\theta})$ for some $\theta\in(0,1)$ and $k$ is known a priori. The set of all possible values of $S$ is denoted by $\mathcal{S}$}, and $S$ is assumed to be uniformly distributed over all $n\choose k$ elements in $\mathcal{S}$}, which is known as the combinatorial prior. We also consider in certain parts of our proofs the i.i.d prior (as an intermediate step), where each item is defective independently with some probability $q\in(0,1)$.
@@
where $\bigvee$ represents the OR operator, $Y_{t}$ is the $t$-th entry of $\mathbf{Y}$, and $\mathbf{X}_{t}=({\mathsf{X}}_{t1},{\mathsf{X}}_{t2},\dots,{\mathsf{X}}_{tn})\in\{0,1\}^{n}$ is the $t$-th row of $\mathsf{X}$}. Given $\mathsf{X}$} and $\mathbf{Y}$, a decoder forms an estimate $\hat{S}$ of the defective set $S$.
@@
While previous works on approximate recovery focused mainly on this symmetric two-sided criterion, our focus in this paper is instead on the following problems, which represent the extremes of placing unequal weighting on false positives and false negatives:
@@
There are natural scenarios where these problems may be more relevant than the previously considered two-sided approximate recovery criterion. For example, SUPERSET may be relevant when trying to contain a highly contagious disease in its initial stages, where a small number of healthy individuals being quarantined is worth ensuring all diseased individuals are identified. An example where SUBSET may be relevant is identifying illegal activity via traitor tracing [26], where it is paramount to avoid false accusations.
@@ I-B Prior Work
Exact recovery. Dorfman’s original scheme [15] was a two-stage adaptive algorithm consisting of splitting up the items into disjoint groups, and including all items of each group in a single test. All items in negative tests are marked as non-defective, while the remaining items in positive tests are subsequently tested individually. While this does indeed beat individual testing, such a scheme is highly suboptimal when the number of defectives is sufficiently sparse relative to the number of items.
@@
In the sparse setting (namely, $k=\Theta(n^{\theta})$ with $\theta\in(0,1)$), there exists a ubiquitous lower bound of $T\geq\big(k\log_{2}\frac{n}{k}\big)(1-o(1))$ on the number of tests required to obtain a vanishing probability of error, regardless of adaptivity or non-adaptivity. This bound is known as the counting bound [4]. The intuition behind this bound stems from the fact that $\log_{2}{n\choose k}=\big(k\log_{2}\frac{n}{k}\big)\big(1+o(1)\big)$ bits are required to describe the defective set, and each test can provide at most 1 bit of information since the test outcomes are binary. This intuition can easily be made precise using standard information theoretic tools such as Fano’s inequality [7, Theorem 1], which implies we cannot obtain an arbitrarily small probability of error with fewer tests. This result has also been strengthened to obtain a strong converse bound stating that $\mathbb{P}(\hat{S}\neq S)\to 1$ as $n\to\infty$ when $T\leq\left(k\log_{2}\frac{n}{k}\right)(1-\varepsilon)$ for arbitrarily small $\varepsilon>0$ [4, Theorem 3.1].
@@
The counting bound can be met when adaptive testing is used for all $\theta\in(0,1)$ using a generalized binary splitting algorithm due to Hwang [21]. The counting bound can also be achieved for $\theta\in\big(0,\frac{\log 2}{1+\log 2}\big)\approx(0,0.409)$ non-adaptively, whereas a converse bound specific to the non-adaptive setting shows that more tests are required for $\theta\geq\frac{\log 2}{1+\log 2}$ [10, Theorem 1.1]. Furthermore, [10, Theorem 1.2] shows that there exists a polynomial-time algorithm matching this bound in the non-adaptive setting. (Prior to their work, the bound was showed to be information-theoretically achievable in [9].)
@@
Approximate recovery. There is also a converse bound specific to the approximate recovery setting [28, Theorem 3], which states that the high probability recovery of a set $\hat{S}$ such that $\max\big\{\lvert\hat{S}\setminus S\rvert,\lvert S\setminus\hat{S}\rvert\big\}\leq\beta k$ for fixed $\beta\in(0,1)$ is not possible when the number of tests satisfies $T\leq(1-\beta)\left(k\log_{2}\frac{n}{k}\right)(1-\varepsilon)$ for arbitrarily small $\varepsilon>0$. A matching upper bound is also known [32], and the polynomial-time algorithm from [10] can be used to achieve the same bound. The techniques in these works appear to be heavily reliant on the approximate recovery criterion being two-sided, and thus not directly suitable for SUBSET and SUPERSET. For instance, in [28] the decoder specifically searches over size-$k$ estimates, which immediately implies an equal number of false positives and false negatives.
@@ I-C Achievable Rates Adapted from Prior Works
With the definitions from Section I-A in mind, we now state the best known existing results for both problems. The relevant existing works stated these results for $\eta^{-},\eta^{+}=\Theta(1)$ in Definition 2, but by slightly adapting their proofs, we produce a stronger result in which these parameters can decay to zero at a certain (sufficiently slow) speed.
@@
Then if $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in[0,1)$, and $\eta^{+}=k^{-o(1)}$, the rate $R^{*}$ is achievable in both the SUBSET and SUPERSET problems.
The condition $\eta^{+}=k^{-o(1)}$ covers a range of possible values, including not only $\eta^{+}=\Theta(1)$ but also the case of decaying sufficiently slowly (e.g., as $(\log k)^{-c}$ for some $c>0$). Moreover, the condition on $\eta^{-}$ allows not only $\eta^{-}=\Theta(1)$ and “slow” decay to zero at rate $k^{-o(1)}$, but also faster (polynomial) decay; by letting $\lambda$ be close to $1$ we can allow $\eta^{-}=\big(\frac{k}{n}\big)^{1-\varepsilon}$ for arbitrarily small $\varepsilon>0$, amounting to at most $k\big(\frac{k}{n}\big)^{1-\varepsilon}$ false negatives in the SUBSET problem. Note also that the term $\zeta(\theta)$ is the optimal exact recovery rate (among all designs) from [10], and is present because achieving exact recovery implies solving both problems even with $\eta^{+},\eta^{-}=0$. This term is dominant for $\theta<\frac{1}{2}$, and for any such $\theta$ the quantity $\eta^{-}k=\Theta(k(\frac{k}{n})^{\lambda})$ is $o(1)$ for $\lambda$ sufficiently close to one, which means that the “approximate recovery” guarantee of SUBSET is in fact equivalent to exact recovery.
For the SUBSET problem, the rate $\log 2\approx 0.693$ is achievable using the Definite Defectives (DD) algorithm along with the near-constant tests-per-item design. See [1] for the algorithm studied under Bernoulli designs, [23] for its study under the improved design, and [2, Section 5.1] for an overview of the extension to approximate recovery.
For the SUPERSET problem, the rate of $\log 2$ is achievable using the Combinatorial Orthogonal Matching Pursuit (COMP) algorithm. See [7] for the algorithm and Bernoulli designs, [23] for its study under the improved test design, and [2, Section 5.1] for an overview of the extension to approximate recovery.
Since the details on these extensions are omitted from [2, Section 5.1], and since the outline therein only covers the case that $\eta^{-},\eta^{+}=\Theta(1)$, we detail the proof of Theorem 1 in Appendix -D.
@@ I-D Contributions
  • We provide a new algorithm for the SUBSET problem that achieves a rate of 1 for all $\theta\in(0,1)$ when $\eta^{-}=\Omega(k^{-\lambda(\frac{1}{\theta}-1)})$ for any $\lambda\in[0,\frac{1}{2})$ (albeit with exponential runtime), thus matching the counting bound and beating the best existing achievability result.
@@ I-E Notation
We generally use boldface letters to denote vectors, lower-case to denote scalars, and capital letters to denote random variables (though the number of tests $T$ is an exception). We use standard Bachmann-Landau notation $o(\cdot),\ \omega(\cdot),\ O(\cdot),\ \Omega(\cdot),\ \Theta(\cdot)$ to denote asymptotic relations. For a matrix $\mathsf{M}$} (resp. vector $\mathbf{v}$) and set ${\mathcal{J}}\subseteq[n]$, ${\mathsf{M}}_{{\mathcal{J}}}$ denotes the sub-matrix of $\mathsf{M}$} (resp. sub-vector of $\mathbf{v}$) corresponding to the columns (resp. entries) indexed by $\mathcal{J}$}. If $\mathcal{J}$$\emptyset$ we use the convention that $\mathsf{M}$$_{\emptyset}$ (resp. $\mathbf{v}_{\emptyset}$) $\emptyset$. For a set $\mathcal{N}$} and some ${\mathcal{U}}\subseteq{\mathcal{N}}$, we denote ${\mathcal{U}}^{\mathsf{c}}\coloneqq{\mathcal{N}}\setminus{\mathcal{U}}$ as the complement of $\mathcal{U}$}. All logarithms have base $e$ unless specified otherwise (e.g., via $\log_{2}$).
@@ II Main Results
While the results in Theorems 2 and 3 establish the optimal rate for SUBSET and SUPERSET in the specified parameter regimes, there are more fine-grained notions of optimality that remain unsolved. In particular, it is unclear whether the range of $\lambda$ (and hence $\eta^{-},\eta^{+}$) in Theorems 2 and 3 can be improved, or if there exists a polynomial-time algorithm for the SUBSET problem that can achieve the same guarantees of that in Theorem 2 (recall that our algorithm has exponential runtime). These questions are further discussed in the following remarks.
Remark 1.
(Discussion on $\eta^{-},\eta^{+}$) While the range of $\lambda$ in Theorems 2 and 3 is not our main focus, we note that the range in Theorem 2 permits $\eta^{-}=\big(\sqrt{\frac{k}{n}}\big)^{1-\varepsilon}$ (i.e., $k\big(\sqrt{\frac{k}{n}}\big)^{1-\varepsilon}$ false negatives) for arbitrarily small $\varepsilon>0$, which is a stronger result than $\eta^{-}=\Theta(1)$ or $\eta^{-}=k^{-o(1)}$. Similarly, according to the range in Theorem 3, when $\theta\in\big(\frac{\log 2}{1+\log 2},\frac{1}{2}\big)$ the converse result holds even when $\eta^{+}=\big(\frac{n}{k^{2}}\big)^{1-\varepsilon}$ (i.e., $k\big(\frac{n}{k^{2}}\big)^{1-\varepsilon}\gg k$ false positives).
While we attempted to find the smallest possible values of $\eta^{-},\eta^{+}$ in Theorems 2 and 3 based on what our own analysis can give, we do not claim these to be the best possible, especially when considering arbitrary non-adaptive designs. In particular, we expect that adopting a near-constant column weight design for SUBSET would lead to smaller values of $\eta^{-}$ being allowed. One way of seeing this is to observe that setting $\theta<\frac{1}{3}$ and $\lambda\approx\frac{1}{2}$ in Theorem 2 permits $\eta^{-}<\frac{1}{k}$, which is equivalent to exact recovery. This is consistent with exact recovery being possible for all $\theta\leq\frac{1}{3}$ under the Bernoulli design. For the near-constant column weight design, this range increases to $\theta\leq\frac{\log 2}{1+\log 2}\approx 0.409$, which suggests that the range of allowed $\lambda$ may be higher, namely $\lambda\in[0,\log 2)$.
Despite being potentially suboptimal in this sense, we stick to the Bernoulli design since it simplifies the analysis, and because making $\eta^{-},\eta^{+}$ as small as possible is not our main focus.
Remark 2.
(Discussion on Computation) The optimal threshold for SUPERSET can be achieved with polynomial-time decoding. When the rate matches that of exact recovery, this follows directly from the results of [10] based on spatial coupling. The only other case is that in which the rate is $\log 2$, which is achieved by the Definitive Defectives (DD) algorithm (Theorem 1). For the SUBSET problem, however, our achievability proof (Theorem 2) is purely information-theoretic and uses a highly inefficient decoder, since the first step of our algorithm uses a decoder from [28] that searches over all ${n\choose k}=e^{\Theta(k\log n)}$ subsets of size $k$, and our second step involves a similar kind of search over $\Omega\big({k\choose 2\eta^{-}k}{n\choose\eta^{-}k}\big)=n^{\omega(1)}$ sets. We leave it as an open problem as to whether the same can be achieved in polynomial time, possibly via the spatial coupling approach of [10]; see Section IV for further discussion.
We can also establish the tightness of these results by comparing them to existing results:
@@
To give some intuition as to why SUPERSET is “harder” than SUBSET, we consider the number of “masked” items, whose status is ambiguous due to only appearing in tests with other defectives (see Definition 6 below for a precise definition):
  • With high probability, for rates above $\log 2$ there are a large number ($\omega(k)$) of masked non-defective items, whereas for rates below $\log 2$ there are relatively few ($o(k)$).
  • With high probability, for rates in $(\zeta(\theta),1)$ (where we recall that $\zeta(\theta)$ is the exact recovery rate), there are some (at least $1$ but $o(k)$) masked defective items, whereas for rates below $\zeta(\theta)$ there are none.
  • For rates above 1, estimates very far from the true defective set become consistent with the test results, so none of the approximate recovery criteria can be attained.
  • SUPERSET is solvable when there are either relatively few masked non-defective items (so they can be guessed as being defective without significantly increasing the size of the estimate), or there are no masked defectives and the rate is below 1 (meaning we can achieve exact recovery and thus solve SUPERSET).
  • In contrast, SUBSET is solvable for rates all the way up to $1$, since the $o(k)$ masked defectives can be guessed as being non-defective without significantly impacting the size of the estimate, and even the presence of $\omega(k)$ masked non-defectives does not impact this.
Note that the above claims on when masking occurs will be evident in our proofs, and can also be inferred from [10].
[…]